OpenStudy (anonymous):

Find the slope of the line tangent to x^3+2xy+y^2=49 at (1,6)

OpenStudy (anonymous):

So find the derivative of the function at 1. Hopefully you know power rules

OpenStudy (anonymous):

Woops just saw the y^2. My bad, totally goofed.

OpenStudy (anonymous):

what do you do with the 2 in 2xy?

OpenStudy (anonymous):

I guess, you simplify to a y= equation first, THEN take the derivative at 1.

OpenStudy (anonymous):

So I believe you get f(x)=-x^3/2-2xy^1/2+49^1/2, then use power rules to find f'(1).

OpenStudy (anonymous):

You know what, I could be completely wrong here. I'm not used to seeing equations in this form and finding derivatives of them, so hopefully a more experienced person will post :P Sorry x.x

OpenStudy (anonymous):

thank you :)

OpenStudy (isaiah.feynman):

Have you tried making it a y= equation?

OpenStudy (isaiah.feynman):

Differentiate the function implicitly.

OpenStudy (raden):

derive implicitly, get x^3+2xy+y^2=49 at (1,6) 3x^2 + 2y + 2xy' + 2yy' = 0 2xy' + 2yy' = -(3x^2 + 2y) y' (2x + 2y) = -(3x^2 + 2y) y' = -(3x^2 + 2y)/(2x + 2y) now subtitute x = 1 and y = 6 into the equation of y' above to get the slope